# Look before Transcription: End-to-End SlideASR with Visually-Anchored Policy Optimization

### v1

@@ B.1 Prompts
Transcribe the speech into text by integrating the speech with the slide content.
### Slide image as context
@@
First, recognize the text in the image and output it within <think> </think>. Then, referring to the thinking content, output the speech recognition result within <answer> </answer>
@@ B.2 Metrics
  • NE-FNR: The false negative ratio of named entities, calculated as $1-\frac{N}{H}$, where $H$ and $N$ denote the recognized and ground-truth entity counts.
@@
B.3 Baselines
Appendix C Results on ChineseLips
Figure 5: Qualitative examples from SlideASR-R.
@@ D.1 Successful Case
Table 7: Ablation results on the SlideSpeech dataset.
ASR OCR VA Dev Test
@@
Table 8: Ablation results on the SlideASR-S dataset.
ASR OCR VA En Zh
@@
Fig 5 shows a comparison of outputs from Qwen2.5-Omni-7B, and Qwen3-Omni-30B-A3B and our proposed VAPO-7B models on samples from the SlideASR-R dataset. Among them, Qwen2.5-Omni-7B uses audio-only input, Qwen3-Omni-30B-A3B uses OCR text extracted from the slide image as context, and VAPO-7B uses the slide image as context input. For Qwen2.5-Omni-7B, due to the lack of auxiliary information, the entity error rate is relatively high. For example, it misrecognized "ConVIRT" as "convert". For Qwen3-Omni-30B-A3B, although slide text is used as context, it fails to utilize it effectively. For example, it also misrecognized "ConVIRT" as "Convert". The VAPO-7B model achieves higher entity recognition accuracy thanks to its "Look before Transcription" reasoning structure.
D.2 Failure Case
Fig 6 reveals a failure mode of our VAPO-7B model, originating from a visual perception error. The issue begins in the <think> block, where the internal OCR component misidentifies a key entity. Specifically, the correct character (pronounced ke) is misrecognized as a different character (pronounced pin). This OCR error is highly plausible because the two characters are visually confusable due to their structural similarity. This type of resemblance is a known challenge for OCR systems, especially with low-resolution text, and it results in the absence of a correct visual anchor.
Table 9: Ablation results of reward function weights on SlideASR-Bench.
Hyperparameter SlideASR-S (en) SlideASR-S (zh) SlideASR-R
$\lambda_{1}:\lambda_{2}:\lambda_{3}:\lambda_{4}$ WER NE-WER NE-FNR WER NE-WER NE-FNR WER NE-FNR
$1:1:1:1$ 4.90 3.19 3.73 2.47 4.21 2.22 27.28 19.31
$1:1:1:2$ 5.27 3.34 3.78 2.50 4.30 2.09 27.67 17.73
$1:1:2:1$ 5.32 4.12 3.91 2.48 4.38 2.09 30.35 22.17
$1:2:1:1$ 5.17 3.45 3.80 2.51 4.23 1.99 27.54 21.18
@@ D.2 Failure Case
Model Setting
Inference time per sample (s)
NE-FNR
@@
Appendix E More Ablation Results
E.1 Ablation Results of Reward Components
Figure 7: More cases of attention visualization.
E.2 Sensitivity Analysis of Reward Weights
To investigate the sensitivity of our VAPO method to the weights ($\lambda_{1}:\lambda_{2}:\lambda_{3}:\lambda_{4}$) of the four reward functions, we conducted an ablation study on the SlideASR-Bench. As shown in Table 9, we compared our default equal weighting scheme (1:1:1:1) against configurations where the weight of the OCR reward ($\lambda_{2}$), ASR reward ($\lambda_{3}$), or the Visual Anchoring reward ($\lambda_{4}$) was doubled.
The results confirm that the 1:1:1:1 configuration provides the best overall performance, achieving the lowest Word Error Rate (WER) across all three subsets. This demonstrates the importance of a balanced approach to synergistic learning.
The analysis also reveals a key trade-off. While doubling the Visual Anchoring reward (1:1:1:2) achieves the best NE-FNR (17.73) on the challenging real-world SlideASR-R set, it does so at the cost of higher overall WER. Conversely, overweighting the ASR reward (1:1:2:1) proved counterproductive, as it discourages the model from leveraging crucial visual context and significantly degrades performance.
Therefore, we conclude that an equal weighting of the reward functions is a simple and effective choice for the VAPO method. This configuration was used for all other experiments reported in this paper.
Appendix F More Cases of Attention Visualization
Figure 6 further presents two cases of attention visualization. It can be seen that when transcribing key entities, the model is able to focus its attention on the same entities in the <think> block. This desirable property enables the model to accurately transcribe key entities in the speech.

### v2

@@ B.1 Prompts
Transcribe the speech into text by integrating the speech with the slide content. ### Slide image as context
@@
First, recognize the text in the image and output it within <think> </think>. Then, referring to the thinking content, output the speech recognition result within <answer> </answer>.
@@ B.2 Metrics
  • NE-FNR: The false negative ratio of named entities, calculated as $1-\frac{H}{N}$, where $H$ and $N$ denote the recognized and ground-truth entity counts.
B.3 Baselines
Appendix C Results on ChineseLips
@@ Appendix D Case Study
Figure 7: Qualitative comparison on the SlideASR-R dataset. We compare the transcription results of the baseline (Qwen2.5-Omni-7B), a strong contextual baseline (Qwen3-Omni-30B-A3B with OCR), and our VAPO-7B. Red text indicates entity recognition errors (e.g., misrecognizing "CheXzero" as "checks zero"), while blue text highlights correctly transcribed entities.
@@ D.1 Successful Case
Fig. 7 shows a comparison of outputs from Qwen2.5-Omni-7B, and Qwen3-Omni-30B-A3B and our proposed VAPO-7B models on samples from the SlideASR-R dataset. Among them, Qwen2.5-Omni-7B uses audio-only input, Qwen3-Omni-30B-A3B uses OCR text extracted from the slide image as context, and VAPO-7B uses the slide image as context input. For Qwen2.5-Omni-7B, due to the lack of auxiliary information, the entity error rate is relatively high. For example, it misrecognized "ConVIRT" as "convert". For Qwen3-Omni-30B-A3B, although slide text is used as context, it fails to utilize it effectively. For example, it also misrecognized "ConVIRT" as "Convert". The VAPO-7B model achieves higher entity recognition accuracy thanks to its "Look before Transcription" reasoning structure.
D.2 Failure Case
Fig. 8 reveals a failure mode of our VAPO-7B model, originating from a visual perception error. The issue begins in the <think> block, where the internal OCR component misidentifies a key entity. Specifically, the correct character (pronounced ke) is misrecognized as a different character (pronounced pin). This OCR error is highly plausible because the two characters are visually confusable due to their structural similarity. This type of resemblance is a known challenge for OCR systems, especially with low-resolution text, and it results in the absence of a correct visual anchor.
Table 9: Ablation results of functions on the SlideSpeech dataset.
ASR OCR VA Dev set Test set
@@ D.2 Failure Case
Table 10: Ablation results of the rewards on the SlideASR-S dataset.
ASR OCR VA SlideASR-S (en) SlideASR-S (zh)
@@
Appendix E More Ablation Results on Reward Functions
Appendix F More Cases of Attention Visualization
Fig. 9 further presents two cases of attention visualization. It can be seen that when transcribing key entities, the model is able to focus its attention on the same entities in the <think> block. This desirable property enables the model to accurately transcribe key entities in the speech.
@@ Appendix F More Cases of Attention Visualization
Model Setting Inference time per sample (s) NE-FNR